An investigator wants to estimate the mean birth weight of infants born full term (approximately 40 weeks gestation) to mothers who are 19 years of age and under. The mean birth weight of infants born full-term to mothers 20 years of age and older is 3,510 grams with a standard deviation of 385 grams. How many women 19 years of age and under must be enrolled in the study to ensure that a 95% confidence interval estimate of the mean birth weight of their infants has a margin of error not exceeding 100 grams? Try to work through the calculation before you look at the answer. 


1
Expert's answer
2021-05-07T09:00:43-0400

"n=(\frac{z_\frac{\alpha}{2 }s} {E}) ^{2 } =(\frac{1.96\times385} {100}) ^{2 } =56.94\approx 57"